# Exact Subgraph Isomorphism Network for Predictive Graph Mining

### v1

@@ 4 Experiments
The first dataset use $H_{p}$ and $H_{n}$, by which positive and negative classes are defined. We first generated a random connected graph, connected with one of $H_{p}$ or $H_{n}$ randomly. The initial random graph has node labels ‘0’, and $H_{p}$ and $H_{n}$ have node labels ‘1’. We generated $300$ instances for each of the positive and the negative classes. We call this dataset ‘Cycle’. See Appendix C.1 for further details of the initial random graph.
@@
As benchmark datasets, we used BZR, COX2, DHFR, and ENZYMES from (Morris et al.,, 2020), and ToxCast and SIDER from (Wu et al.,, 2018).
@@
For performance comparison, we used GNN methods such as GCN (Kipf and Welling,, 2017), GAT (Velickovic et al.,, 2018), GATv2 (Brody et al.,, 2022), GIN (Xu et al.,, 2019), PNA (Corso et al.,, 2020), GNN-AK (Zhao et al.,, 2022), and PPGN (Maron et al., 2019a, ). For node attributes, EIN only used discrete node labels, while compared GNN methods also incorporate continuous node attributes. GCN, GAT, GATv2, and GIN optimizes the number of the units $\{64,128,256\}$ and the number of epoch by the validation set, and other settings follow (You et al.,, 2021). PNA optimizes the number of the units $\{16,32,64\}$ and the number of epochs by the validation set, and the number of the message passing is set as $2$. GNN-AK uses GIN as the base GNN. Other settings of GNN-AK and PPGN follow the author implementation. For EIN, $\text{maxpat}\in\{5,10\}$, $K\in\{2,6,10\}$, and $\sigma\in\{\mathrm{sigmoid},\mathrm{LeakyReLU}\}$ are selected by the validation performance. We also evaluate performance of the combination of EIN and GNN described in § 2.4. For GNN, we employed GIN and the combine model is denoted as EIN+GIN. For other details, see Appendix C.2.

### v2

@@ 4 Experiments
The first dataset use $H_{p}$ and $H_{n}$, by which positive and negative classes are defined. We first generated a random connected graph, connected with one of $H_{p}$ or $H_{n}$ randomly. The initial random graph has node labels ‘0’, and $H_{p}$ and $H_{n}$ have node labels ‘1’. We generated $300$ instances for each of the positive and the negative classes. We call this dataset ‘Cycle’. See Appendix A.1 for further details of the initial random graph.
@@
As benchmark datasets, we used BZR, COX2, DHFR, ENZYMES, and PTC MR from (Morris et al.,, 2020), and ToxCast from (Wu et al.,, 2018).
@@
For performance comparison, we used GNN methods such as GCN (Kipf and Welling,, 2017), GAT (Velickovic et al.,, 2018), GATv2 (Brody et al.,, 2022), GIN (Xu et al.,, 2019), PNA (Corso et al.,, 2020), GIN-AK (Zhao et al.,, 2022), PPGN (Maron et al., 2019a, ), and ESC-GNN (Yan et al.,, 2024). For node attributes, EIN only used discrete node labels, while compared GNN methods also incorporate continuous node attributes. GCN, GAT, GATv2, and GIN optimizes the number of the units $\{64,128,256\}$ and the number of epoch by the validation set, and other settings follow (You et al.,, 2021). PNA optimizes the number of the units $\{16,32,64\}$ and the number of epochs by the validation set, and the number of the message passing is set as $2$. Other settings of GIN-AK, PPGN, and ESC-GNN follow the author implementation. For EIN, $\text{maxpat}\in\{5,10\}$, $K\in\{2,6,10\}$, and $\tau_{\rm max}\in\{1,30\}$ for $\boldsymbol{\theta}$ ($\tau_{\rm max}$ for $\boldsymbol{b}$ is fixed as $1$) are selected by the validation performance. We also evaluate performance of the combination of EIN and GNN described in § 2.4, in which GIN is employed as the combined GNN. In the results, the single EIN is denoted as EIN-$L_{0,2}$ and the EIN combined with GIN is denoted as EIN-$L_{0,2}$-GNN. For both of EIN-$L_{0,2}$ and EIN-$L_{0,2}$-GNN, the candidate of the sparsity $s$ is ${\mathcal{S}}=\{1,5,10,25,50,75,100\}$. For other details, see Appendix A.2.